Phantomdex:

A principal amount of $2,500 is placed in a savings account with an APR of 4.1% compounded semi-annually for 7 years. The same principal amount is placed in a savings account with an APR of 2.2% compounded continuously for 10 years. Which account earns more interest? The semi-annually compounded interest account earns more, with $615.19 in total interest earned. The continuously compounded interest account earns more, with $821.41 in total interest earned. The semi-annually compounded interest account earns more, with $821.41 in total interest earned. The continuously compounded interest account earns more, with $615.19 in total interest earned.

Arizona:

ok did u use the site that the other user gave u to help me

Arizona:

oops i met You

Phantomdex:

I'm trying but it says apr not interest rate..unless their the same thing

Arizona:

hmm @axie can u help again

Phantomdex:

@arizona wrote:
hmm @axie can u help again
I don't think they will cause they gave me the google answer for the last one sobs

dontsaymyname:

Hold Please

dontsaymyname:

alrighty, are you comfortable using a calculator to solve?

Phantomdex:

I believe so

dontsaymyname:

\[2500 (1.02) ^14\]

dontsaymyname:

sorry the 14 is weird format, but its squared by 14 i believe

Phantomdex:

Oh that's fine, my calculator just doesn't seem to have a squaring thing on it

dontsaymyname:

okay let me see what i can do on my end

dontsaymyname:

Okay it seems, 2500 (1+interest semi(meaning 2)annually(year)) Squared by years

dontsaymyname:

soo 2500 (1.02) Squared by 14 because the 7 would be annually and we have semi so 7*14, equalling 3298.697

dontsaymyname:

but they want us to deicde which makes more, so let me try the 2nd one

SmoothCriminal:

To compare the two savings accounts, we can utilize the formula for compound interest: For the semi-annually compounded account: \[ A_{\text{semi-annual}} = P \left(1 + \frac{r}{n}\right)^{nt} \] Where: \(A_{\text{semi-annual}}\) = the future value of the investment/loan, including interest \(P\) = the principal investment amount (the initial deposit or loan amount) = $2,500 in both cases \(r\) = the annual interest rate (decimal) = 4.1% or 0.041 for the semi-annually compounded \(n\) = the number of times that interest is compounded per year = 2 for semi-annually compounded \(t\) = the time the mazuma is invested or borrowed for, in years = 7 years for the semi-annually compounded Now, let's calculate the future values for both accounts: \[ A_{\text{semi-annual}} = 2,500 \left(1 + \frac{0.041}{2}\right)^{2 \cdot 7} \approx \$2,974.27 \] For the perpetually compounded account: \[ A_{\text{continuous}} = 2,500 \cdot e^{0.022 \cdot 10} \approx \$3,161.84 \] Now, let's calculate the interest earned for both accounts: \[ \text{Interest}_{\text{semi-annual}} = A_{\text{semi-annual}} - P = \$2,974.27 - \$2,500 = \$474.27 \] \[ \text{Interest}_{\text{continuous}} = A_{\text{continuous}} - P = \$3,161.84 - \$2,500 = \$661.84 \] So, the perpetually compounded interest account earns more with \$661.84 in total interest earned.

dontsaymyname:

Well i guess try that LOL im sorry

SmoothCriminal:

@dontsaymyname wrote:
Well i guess try that LOL im sorry
JJ did amazingly 🙏

dontsaymyname:

I tried but I'm glad u were able to show the work

Phantomdex:

I have more questions 0-0

dontsaymyname:

OH DEAR

SmoothCriminal:

@phantomdex wrote:
I have more questions 0-0
Sure

dontsaymyname:

i found a worksheet that shows it by hand but i was tryna calculate what the heck "e" was LOL

Phantomdex:

fair anyways onto the next

Kylah2023:

@kingsteve